If 750 ml of a 6% solution is diluted to 2000 ml, the percentage of the final solution will be?
Answer by jim_thompson5910(35256)

When you dilute any solution, you're just adding pure water to it (to lower the concentration of the pure stuff in relation to the mix).


We have 750 ml of a 6% solution, so we have 750*0.06 = 45 ml of pure stuff (whatever it is).


Notice how (Amount of pure stuff)/(Amount of solution) = 45/750 = 0.06 which is 6% as expected.


----------------------------------------------------------------------------


We bump up the total amount of solution to 2000 ml (2000-750 = 1250 ml of pure water is added). The amount of pure stuff (whatever it is) stays the same at 45 ml. We have this new fraction


(amount of pure stuff)/(amount of solution) = 45/2000 = 0.0225


Multiply by 100 to convert over to a percentage: 0.0225*100 = 2.25%


So the percentage of the final solution is 2.25%